### v1

# Machine Learners Should Acknowledge the Legal Implications of Large Language Models as Personal Data

Abstract
Does GPT know you? The answer depends on your level of public recognition; however, if your information was available on a website, the answer is probably yes. All Large Language Models (LLMs) memorize training data to some extent. If an LLM training corpus includes personal data, it also memorizes personal data. Developing an LLM typically involves processing personal data, which falls directly within the scope of data protection laws. If a person is identified or identifiable, the implications are far-reaching: the AI system is subject to EU General Data Protection Regulation requirements even after the training phase is concluded. To back our arguments: (1.) We reiterate that LLMs output training data at inference time, be it verbatim or in generalized form. (2.) We show that some LLMs can thus be considered personal data on their own. This triggers a cascade of data protection implications such as data subject rights, including rights to access, rectification, or erasure. These rights extend to the information embedded with-in the AI model. (3.) This paper argues that machine learning researchers must acknowledge the legal implications of LLMs as personal data throughout the full ML development lifecycle, from data collection and curation to model provision on, e.g., GitHub or Hugging Face. (4.) We propose different ways for the ML research community to deal with these legal implications. Our paper serves as a starting point for improving the alignment between data protection law and the technical capabilities of LLMs. Our findings underscore the need for more interaction between the legal domain and the ML community.
Machine Learning, ICML
1 Introduction
GPT probably knows you—not just you, but it has likely memorized data about countless other individuals as well (Verma et al., 2024). Most Large Language Models (LLMs) memorize parts of the input data provided during training in verbatim or quasi-verbatim manner (Carlini et al., 2022; Somepalli et al., 2023; Verma et al., 2024). This memorization is double-edged. The training objective of an LLM is partly generalization, but memorization is also an essential component (Tirumala et al., 2022; Power et al., 2022; Hartmann et al., 2023; Biderman et al., 2024). Without memorization, an LLM would be, for example, unable to tell us that the Eiffel Tower is located in Paris. Thus, memorization is an important and — to some extent — deliberate aspect of training LLMs. While memorization is a crucial feature of LLMs, it also comes with challenges (Bender et al., 2021), while a key problematic aspect is the consideration of memorized personal data.
Although the problematic nature of personal data memorization in LLMs has been mentioned numerous times (Carlini et al., 2019; Hu et al., 2022; Carlini et al., 2022; Zhang et al., 2023a; Cooper & Grimmelmann, 2024), the debate about its implications is still in its early stages. While much of the technical work focuses on unlearning techniques (Juliussen et al., 2023; Bourtoule et al., 2021; Zhang et al., 2023b), the legal implications are often overlooked in the technical ML community [^1] We use the term “ML community” to refer to researchers — whether from academia or industry — who work on topics suitable for publication at conferences like ICML, NeurIPS, or ICLR.. For example, the processing of personal data requires a legal basis. Furthermore, individuals whose data has been memorized by an LLM may have the right to request its erasure from the model, which could require the removal of that information from the model itself, see discussion (Villaronga et al., 2018; Yaish, 2019; Fabbrini & Celeste, 2020; Allegri et al., 2022; Juliussen et al., 2023). These questions are already unfolding in real life. For instance, complaints have been filed against OpenAI, alleging General Data Protection Regulation (GDPR) [^2] EU Regulation 2016/679, 27.4.2016. violations due to the company’s failure to ensure the accuracy of personal data, comply with access and rectification requests, and provide transparency about the sources and processing of personal data in ChatGPT (NOBY, 4). From our perspective, not all model developers fully consider or acknowledge the serious implications of this requirement.
Why should ML researchers care whether LLMs are personal data or not? If LLMs qualify as personal data, ML researchers may be responsible for their processing. In such cases, researchers must also comply with data protection regulations. As we demonstrate below, non-compliance with data protection regulations can lead to severe fines, whether on a personal, university, or company level. Please also note that some data protection laws, such as the GDPR, even extend to researchers outside the EU as long they process data from people within the EU (see Article 3 GDPR). Beyond legal compliance, successful LLM research (and its applications) requires not only technical rigor but also the trust of users. We believe that adhering to data protection regulations can enhance trust in ML systems. When researchers commit to data protection safeguards, their applications are more likely to gain public acceptance.
Computational Perspective. LLMs clearly memorize some input data verbatim in normal usage contexts while still generalizing to new contexts (Carlini et al., 2019; Hartmann et al., 2023; Tirumala et al., 2022; Biderman et al., 2024). This memorization increases with the growing capacity of the model, the frequency of examples in the training data, and the number of tokens of context used to prompt the model (Carlini et al., 2022). If data are not outputted verbatim, in some settings, an adversarial attacker is able to extract large amounts of training data (Nasr et al., 2023a). However, it has also been noted that verbatim memorization of facts is an important feature — and not a bug — of LLMs (Biderman et al., 2024). A weaker form of verbatim memorization is the knowledge of facts in LLMs (Ippolito et al., 2022). LLMs store certain facts (Petroni et al., 2019; Hu et al., 2023; Roberts et al., 2020; Yu et al., 2022; Wang et al., 2023) which also include personal data (Huang et al., 2022; Kim et al., 2024). But why do some LLMs store personal data? Training LLMs relies on large web-scraped datasets such as Common Crawl (Crawl, 2008) or The Pile (Gao et al., 2020), see also (Villalobos et al., 2024) for a general discussion. Since the internet contains a lot of personal data, this information becomes part of the datasets as well. Memorization of personal data also scales with the capacity of the model (Lu et al., 2024). While on the one hand the memorization is not perfect (Elazar et al., 2021; Cao et al., 2021), it is possible to alter specific facts in LLMs (Meng et al., 2022; Peng et al., 2023).
Legal Perspective. The relationship between AI memorization and data protection law under the GDPR has garnered considerable attention (Wachter et al., 2017; Hacker, 2021; Hacker et al., 2023; Cooper & Grimmelmann, 2024). For this reason, we focus on the GDPR in this paper. While issues of memorization are also highly relevant in the context of copyright law–as highlighted by the New York Times’ lawsuit against Microsoft and OpenAI (Cooper & Grimmelmann, 2024; Sag, 2023; Pope, 2024)–the data protection implications of memorization are the primary focus of this paper. Since LLMs are commonly trained on publicly available internet data, they inevitably pick up personal information. Notably, personal data does not lose its legal status simply because it has been made public. Thus, if this training data is memorized by a model through training, the model itself could potentially be considered personal data under the GDPR (Veale et al., 2018; Juliussen et al., 2023; EDPB, 2024) and trigger unexpected legal consequences. This debate has gained international traction with both the Hamburg Data Protection Commissioner and the Danish Data Protection Authority asserting that LLMs do not contain personal data (Hamburg, 2024; Datatilsynet, 2023), arguing that the data is transformed into abstract mathematical representations and probability weights. However, this position has sparked considerable controversy (Coyer, 2024). The European Data Protection Board [^3] The European Data Protection Board is an independent EU body that ensures the consistent application of the GDPR and provides related guidance, see https://www.edpb.europa.eu/about-edpb/who-we-are/european-data-protection-board_en for more details . has joined the discussion, issuing an opinion stating that AI models trained on personal data cannot automatically be considered anonymous, implying that in most cases, they must be classified as personal data. From our perspective, this stance seems to stem from policy considerations that do not align with the current legal landscape. Instead, it appears to be driven by concerns over the impracticality or undesirability of the legal implications under data protection law. This conflation of technical facts and legal implications has led to misleading conclusions that fail to address the nuanced complexities of ML systems.
ML Research should acknowledge legal implications of LLMs as personal data. The current literature lacks a comprehensive paper that consolidates the various opinions and arguments as to whether LLMs [^4] Please note that in our work, as in the previous (legal) literature, we focus on LLMs. However, our arguments also apply to other modalities, such as Vision Language Models. qualify as personal data under the GDPR. Furthermore, from our point of view, legal implications of this classification remain underexplored, and many developers and computer scientists are unaware of the severe legal implications this issue entails. Therefore, the goal of this paper is twofold: first, we aim to clarify the legal status of LLMs trained on personal data under the GDPR. Second, we argue that ML researchers must acknowledge the legal implications of LLMs as personal data and consider these circumstances during model development. To facilitate this understanding, we contribute the following:
  • Section 2 gives a primer on data protection under the GDPR for ML scientists.
  • Section 3 reiterates that LLMs may memorize personal data and, if this data can be extracted, the LLMs themselves must be treated as personal data.
  • Section 4 highlights that the legal implications of memorization are severe and currently not recognized by ML scholars
  • Section 5 concludes by proposing research directions to help ML scholars address these legal challenges effectively.
2 EU Data Protection Law: A Primer for Computer Scientists
This section provides ML researchers with a basic understanding of data protection law under the GDPR. We start with a general introduction to the GDPR framework and then explain what qualifies as personal data and the legal implications of processing it.
2.1 Personal Data and Responsibilities of Stakeholders
The GDPR forms the central legal framework for the protection of individuals with regard to the processing of their personal data within the European Union (Jones & Kaminski, 2020). According to Article 4(1) GDPR, personal data is defined as “any information relating to an identified or identifiable natural person.” In other words, if information can be linked back to a person, whether directly or indirectly, it is protected under the GDPR (Finck & Pallas, 2020). Typical examples that first come to mind for personal data in the context of ML applications are names or dates of birth. However, the Court of Justice of the European Union (CJEU) clarified that personal data is “not restricted to information that is sensitive or private, but potentially encompasses all kinds of information, not only objective but also subjective.” [^5] Case C-434/16 Nowak [2017] ECLI:EU:C:2017:994, para 34. This interpretation underscores the broad scope of personal data (Purtova, 2018) and demonstrates that even information that may not initially appear to relate to an individual can still qualify as personal data. For instance, even parameters within a model that encode seemingly anonymized patterns—such as weights linked to frequent phrases in text or common visual features—may still be considered personal data.
The GDPR distinguishes between two key roles: the data subject and the data controller. The data subject is the natural person whose personal data is processed, [^6] Article 4(1) GDPR. e.g., the person whose information is stored in the LLM. The data controller, on the other hand, is the entity that determines the purposes and means of processing the data. [^7] Article 4(7) GDPR. In ML research, determining the data controller is not straightforward. Could it be the principal investigator, the university, or even the ML researcher? Typically, organizations such as companies or institutions act as data controllers because they define how and whether personal data should be processed by their employees. [^8] Case C-131/12 Google Spain and Google [2014] ECLI:EU:C:2014:317, para. 33. They have significant authority over data processing decisions, which qualifies them as data controllers under the GDPR. In the context of ML research, however, the situation is different. Researchers generally enjoy considerable independence due to their constitutionally protected freedom of science and teaching. [^9] See for EU researchers: Article 13 Charter of Fundamental Rights of the European Union. This independence allows them to make their own decisions about the purposes, sources, and methods of processing personal data. Whoever autonomously decides, for example, which datasets to use for training or how to pre-process the data, may qualify as a data controller under the GDPR. Therefore, it has been argued that researchers or principal investigators can be data controllers as well (Research & Innovation, 2024).
While the data subject holds rights, the data controller is bound by obligations. The GDPR grants several rights to data subjects, such as the right to access their data, to request its erasure (the “right to be forgotten”), to object to its processing, and to have their data provided in a portable, machine-readable format (Vrabec, 2021; Wolters, 2018). Conversely, data controllers must ensure data confidentiality, notify authorities of breaches, conduct risk assessments, and process data only when there is a lawful basis, such as consent, and for a specific purpose and limited duration (Hintze, 2018).
2.2 When Does Data Relate to an Individual?
Under the GDPR, information is considered personal data if it relates to an identified or identifiable person. A person is “identified” when their identity can be directly determined from the information itself. [^10] Case C-582/14 Breyer [2016] ECLI:EU:C:2016:779, para 38. For example, details like a name, date of birth, address, or medical history clearly identify the individual if memorized in an LLM. Moreover, a person is considered “identifiable” if–while the information alone is not sufficient to determine their identity–it becomes possible to do so when combined with additional data (Article 4 (2) GDPR). For example, a dataset that contains browser histories without names might still be personal data. When combined with IP addresses or patterns of visited sites, it can reveal an individual’s identity. Again, if identifiable data are memorized in an LLM, the model itself must be seen as personal data. To determine whether someone is identifiable, the GDPR requires considering all means likely to be used by the data controller or any other party to identify the individual (Recital (26) GDPR), see also Finck & Pallas (2020) for further discussion. This assessment must include factors such as the cost, time, and resources needed for identification, as well as the current state of technology and potential future technological advancements (Finck & Pallas, 2020).
2.3 The GDPR Trigger: Data Processing
Many rights and obligations of the GDPR are triggered whenever personal data is “processed”. “Processing” is defined as any operation performed on personal data (see Article 4(2) GDPR). Article 4(2) GDPR lists examples of processing activities, including collection, organization, structuring, storage, retrieval, use, disclosure, erasure, or destruction. If personal data can be retrieved from an LLM that memorized it during training within its parameters, the model itself should be classified as personal data. Consequently, any action performed on the model, such as training, uploading, downloading, storing, or deploying it on platforms like GitHub or HuggingFace, fine-tuning, or otherwise sharing it, constitutes processing of personal data.
3 The Implications of Memory in LLMs
ML developers who act as data controllers under the GDPR (see Section 2.1) must comply with its provisions whenever personal data is used for training. But what happens after training is complete? Does the LLM memorize personal data in a way that it remains “stored” within the model? If so, this raises the critical question of whether a fully trained model should itself be classified as personal data and therefore fall within the scope of the GDPR (EDPB, 2024; Veale et al., 2018; Leiser & Dechesne, 2020; Juliussen et al., 2023). This would trigger significant implications and challenges that can be difficult to comply with in practice (see Section 4). In this section, we briefly discuss both questions and illustrate why LLMs qualify as personal data.
3.1 Why LLMs Memorize Personal Data
Figure 1: LLMs and partially successful personal data editing techniques. (a) We aim to edit facts about personal data, in our case the president of the US, in GPT-J 6B with MEMIT (Meng et al., 2022, Mass-Editing Memory in a Transformer). (b) Pre-interventional model. The heatmap shows the influence of each word on the prediction across all layers. The test prompt yields the correct answer. We also include the response of an unrelated and related prompt to our intervention. Parentheses include the likelihood of the response. (c) Edited model. The test prompts yield the desired answer, the unrelated prompt remains unchanged, and the intervention-related prompt, however, still gives the unedited answer. The latter highlights the challenges of editing facts and data removal in LLMs. Icons from flaticon.com
LLMs store training data in parameters through an interconnected and overlapping architecture. This complex structure makes it difficult to directly access or prove data memorization. But does personal data have to be human-perceptible? Under the GDPR, the format in which information is encoded is irrelevant when determining whether it qualifies as personal data (EDPB, 2024). For instance, JPEG, MP3, or PDF files are personal data if they allow an identification of an individual, in spite of their contents not being directly perceptible. Similarly, the fact that data is imperceptibly stored within AI model parameters does not preclude it from being considered personal data.  Furthermore, pseudonymization does not alter the status of personal data. Personal data that cannot be attributed to a specific individual without additional information is still regarded as personal data (Recital 26 GDPR). Personal data stored in AI models can be considered a form of pseudonymization (Veale et al., 2018).  Similarly, the imperceptibility of data stored in LLMs parameters does not exclude the possibility for personal data to be stored (EDPB, 2024). Moreover, just because we cannot directly access, observe, or pinpoint specific data within the model, this does not mean that it is not stored. To date, evidence of stored data in LLMs can only be inferred by observing correlations between inputs and outputs (Cooper & Grimmelmann, 2024). Studies estimate that LLMs can memorize 0.1 to 10 percent of their training data verbatim (Cooper & Grimmelmann, 2024). In addition, evidence suggests that larger models memorize more than smaller ones, and that data that is frequently repeated in the training set is more likely to be memorized (Lee et al., 2022; Carlini et al., 2023; Nasr et al., 2023b).
Figure 1 shows a real-world example of GPT-J 6B storing personal data and model attribution. We used the tools from Meng et al. (2022) to edit the model. GPT-J 6B predicts the next word with the input “The last name of the president of the USA is” with the correct sequence “Trump”. Please note that we choose personal data that could also be considered common knowledge on purpose. We do not want to expose the personal data of a less-known individual. Our argument works with any personal data, whether the person is of public interest or not. The heatmap in Figure 1b also shows the attribution of different layers and the input sequence to the output prediction. Editing personal data will be discussed in Section 5.
Some algorithms, like k-nearest neighbor classification and support vector machines, explicitly encode data points and make (personal) training data an integral part of the model (Brown et al., 2021). However, even neural networks have to balance memorization and generalization. In fact, research suggests some degree of memorization is essential for generalization (van den Burg & Williams, 2021; Chatterjee, 2018; Feldman & Zhang, 2020). Furthermore, compression through training reduces data but preserves essential details, much like a ZIP file (Cooper & Grimmelmann, 2024). As a result, even highly compressed models are likely to retain identifiable patterns or characteristics.
3.2 When Do LLMs Qualify as Personal Data?
Determining whether LLMs qualify as personal data under the GDPR depends on whether it is “reasonably likely” that an individual can be identified through the model’s output. This requires a case-by-case assessment and depends on the effort needed to retrieve personal data in the model (EDPB, 2024). A LLM can only be considered non-personal data, if no personal data from the training data can be extracted using reasonable means; and its outputs do not relate to the natural persons whose data was used to train it. This requires an impossibility to single out, link, and infer information from the supposedly anonymous LLM (EDPB, 2024)—a very high bar to clear.
3.3 Alternative Views
Our position is challenged by scholars and data protection authorities who argue that AI models, in general, cannot be classified as personal data (Wachter & Mittelstadt, 2019; Leiser & Dechesne, 2020; Datatilsynet, 2023; Hamburg, 2024). Consequently, they contend that the legal implications outlined in Section 4 do not apply. Their argument is that even specific AI models, such as LLMs, do not memorize data. Therefore, the argument goes, LLMs cannot, on their own, be considered personal data under the GDPR. Specifically, LLMs are compared to statistical reports, which are not regarded as personal data if they contain only conclusions and aggregated data derived from statistical analysis (Datatilsynet, 2023). It is further argued that personal data used to train LLMs is transformed into abstract mathematical representations and probability weights, which supposedly is not personal data (Hamburg, 2024). Instead of memorizing personal data, LLMs are said to merely learn correlations between tokens based on probability weights (Hamburg, 2024).
However, these arguments are flawed for the following four reasons: First, the format of the encoded information is irrelevant under the GDPR (see Section 1). Second, unlike a statistical report, an LLM can be queried, for instance, with an individual’s name to produce personal data. Third, while it is undisputed that the outputs of LLMs are probabilistic [^11] Even setting the “temperature” parameter of an LLM to zero does not lead to complete determinism (Ouyang et al., 2023). and often inaccurate (see Figure 2), data does not need to be fully accurate to be personal data under the GDPR. Even blurred images or partially incorrect details may suffice for identification, as fragments or partial replicas can reveal an individual’s identity depending on the context. Fourth, our experiments (see Figure 1) demonstrate that personal data is stored within LLMs and can even be edited. As a result, if personal data stored within an LLM can be extracted in a way that makes the identification of an individual “reasonably likely”, the LLM must be considered personal data under the GDPR (EDPB, 2024; Veale et al., 2018; Juliussen et al., 2023).
4 Overlooked Legal Implications of LLMs as Personal Data by the ML Community
Classifying an LLM as personal data carries profound legal implications for the machine learning community, potentially impacting data handling practices, compliance requirements, and model development processes. In the following, we will discuss different non-exhaustive legal requirements: First, researchers need a legal basis if they are training a model or providing the model on, e.g., GitHub or Hugging Face, such as consent, contract, or legitimate interest. Second, data subjects would have a right to access, delete, and rectify their personal data within the LLM. In the following, we focus on the regulations regarding a lawful basis, the right of access and deletion, and the obligation of data controllers to ensure privacy by design since these are, from our perspective, the most overlooked requirements in ML. Please note that we simplify some of the legal nuances for for clarity and due to space limitations, see (Feiler et al., 2018) for legal details.
Figure 2: Large Language Models are always probabilistic. (a) When GPT-4o is asked to choose a color, we observe probabilistic behavior. The y-axis shows the probabilities output by the OpenAI API. (b) When asked who was the president of the USA in 2023, the probabilistic behavior is hidden by a high-likelihood single response.
Article 89 GDPR states that the processing of personal data for scientific purposes shall be subject to appropriate safeguards. One of these safeguards is, for example, the principle of data minimization. From our point of view, this is hardly seen in LLM research; in contrast, more and more data are used (Bender et al., 2021; Shi et al., 2023; Kaplan et al., 2020). Article 89 of the GDPR also allows EU member states to provide derogations in the context of scientific research with regard to the rights of data subjects. However, the scope of Art. 89 is highly controversial (See (Kindt et al., 2021; Biega & Finck, 2021) for more details). Because of this controversial debate, we focus on the EU-wide law.
4.1 Lawful Basis: Researchers Require a Reason to Process Data
Whenever an LLM is classified as personal data, any “processing” of the model requires a legal basis (see Section 2.3 for an explanation of the term processing; A full list of legal bases can be found in Article 6 GDPR). Therefore, whenever a model is shared, uploaded, downloaded, or distributed locally or on platforms such as GitHub or Hugging Face, these processes require a legal basis. Under the GDPR, the main legal bases for data processing include consent, the performance of a contract, protection of vital interests, and legitimate interests (Article 6(1) GDPR). These legal bases are complex and raise a number of legal issues, which we will not address here. Instead, we refer the interested reader to (Feiler et al., 2018).
4.2 Right of Access: Researchers Have to Provide Individuals With Information About Their Personal Data Stored in LLMs
According to Article 15 GDPR, data subjects have a right to obtain access to their personal data from the controller. This raises the question of how this right can be implemented, as it is almost impossible to immediately access specific data within the LLM, as mentioned above. It is unclear how data controllers can comply with such a request. They could either use structured prompts to query the model for potentially relevant personal data (Carlini et al., 2021), or ask data subjects to suggest their own prompts. Testing prompts under different conditions, similar to red teaming (Perez et al., 2022), and parameters can help identify data that is consistently included in the outputs, indicating a higher likelihood of personal data being stored. In addition, the information provided to the data subject would need to include a clarification that a complete extraction of all personal data from the model is not technically feasible, that the model may hallucinate, and that any answers should be presented as approximations.
4.3 Right to Be Forgotten: At the Request of Data Subjects, Personal Data Must Be Deleted From LLMs
Furthermore, data subjects have, under some circumstances, the right to request the erasure of their personal data from the controller within an LLM (Article 17 GDPR). However, fulfilling this obligation poses technical challenges (Veale et al., 2018; Villaronga et al., 2018; Cooper et al., 2024). Simply retraining the model is often not possible, keeping in mind that the pure energy costs for training an LLM can be several million dollars (de Vries, 2023). Different approaches to solve these issues have been discussed in ML Literature (Veale et al., 2018; Villaronga et al., 2018). Approaches for fast and easy “machine unlearning” (Bourtoule et al., 2021) have only recently been proposed and are still largely unexplored, let alone ready for use(Nguyen et al., 2022; Zhang et al., 2023b). The methods currently under discussion cannot be retrofitted into existing systems but would require a complete redesign of the entire model pipeline with unclear implications.
4.4 Data Protection by Design: Researchers Are Required to Think About Data Protection From the Beginning
In all cases, LLM developers must implement technical and organizational measures to comply with data protection principles (Article 25(1) GDPR). For example, developers must make their models resistant to LLM-specific privacy attacks (Yao et al., 2024; Li et al., 2023), such as data extraction attacks or membership inference attacks. Furthermore, Article 35(1) GDPR requires a data protection impact assessment for specific risky areas. Machine Learning, and specifically LLMs qualify as such a risky area (European Data Protection Supervisor, 2019). Developing this data protection impact assessment is a non-trivial task (Kloza et al., 2019). It usually requires critical thinking, extensive documentation and providing safeguards in advance.
4.5 Possible Legal Consequences for LLM Researchers
The legal consequences of classifying an LLM as personal data under the GDPR are far-reaching. For example, if data controllers transfer a model without a legal basis to do so, and such a model is inverted, this would likely be considered both a data breach and a violation of the principle of security in general (Veale et al., 2018). In addition, breaches of obligations and rights of data subjects also trigger significant consequences. For particularly serious breaches, as defined in Art. 83(5) GDPR, the fine can be up to 20 million euros or, in the case of a company, up to 4 percent of its total worldwide turnover in the previous financial year, whichever is higher. But even the catalog of less serious breaches in Art. 83(4) GDPR provides for fines of up to 10 million euros or, in the case of a company, up to 2 percent of its total worldwide turnover in the previous financial year, whichever is higher. The fines are not merely a theoretical risk, as there have already been cases where individual researchers have faced penalties (APD, 2022) as well as at least 35 fines against universities (GDPR ET, 2024), see also (Ruohonen & Hjerppe, 2022) for more details.
4.6 Summary of Legal Implications of LLMs as Personal Data
LLMs can include personal data, which has significant legal implications for the ML community that need to be acknowledged. First, ML Researchers need a legal basis to process the data. Second, this includes not only training the model itself but also the distribution of the models on platforms like GitHub or Hugging Face. Furthermore, individuals have a right of access and a right to be forgotten. Before processing begins, ML researchers need to make a data protection impact assessment to ensure privacy by design.
5 Recommendation for Research Practice in the LLM Community
In this section, we provide brief recommendations for ML researchers in the LLM domain.
The GDPR applies to the processing of personal data. If a model is trained without any personal data, it falls outside the scope of the GDPR. [^12] Note that in some rare cases, an LLM’s input data can generate personal data as output, even if that data was not stored in the model but was generated due to the input. For example, if a user asks the model to spellcheck a CV, the model will output personal data. The legal implications of these exceptional cases are beyond the scope of this paper. In this case, the GDPR does not apply. Researchers who train LLMs with personal data need to ensure that they have a legal Data used for training may require a legal basis or fall under a specific scientific exception. Even in the latter case, other obligations may still apply, such as the obligation for data minimization (Article 89(1) GDPR). It might also be possible to use Differential Privacy methods (Dwork, 2006) to avoid access of personal data by users of the model (Cummings & Desai, 2018; Holzel, 2019), see (Cummings et al., 2023) for more details.
To prevent the exposure of personal data stored in an LLM, the model could be encapsulated in a privacy-preserving framework. Additional layers can be implemented to minimize the risk of disclosing stored personal data (Cooper & Grimmelmann, 2024). For example, at the “front end”, user input can be filtered or modified before being processed by the model. On the back end, models can be tuned to refuse to generate content that could lead to the re-identification of individuals. Finally, again on the “front end”, LLM outputs can be filtered or modified before being delivered to users. This layered architecture allows developers to mitigate the risk of exposing stored personal data. Consequently, it could be that some providers of LLMs may not need to remove specific data points from their trained model, but filter results at the output level.
On a technical side, it might be possible to use and extend unlearning methods (Bourtoule et al., 2021; Nguyen et al., 2022) to comply with the right to be forgotten. The idea of machine unlearning is that it is possible to let the model “forget” specific data points. It might also be possible to use editing methods like MEMIT (Meng et al., 2022), which we used in the context of GPT-J 6B in Figure 1 to edit facts of personal data and thus remove them from the data. In Figure 1b, we edited the last name of the president of the USA from Trump to the president of France, “Macron”. After editing, the model outputs the correct information for the initial prompt. However, while unlearning tools in the context of LLMs pose a promising direction, there are still severe challenges (Xu et al., 2024). This also shows our example in Figure 1b. When using the “last name of the president of the USA”-unrelated prompt about the Eiffel Tower, the model answers unchanged. In contrast, asking it in a different way about the president of the US still results in the model memorizing the correct answer of “Trump”.
Lastly, one might wonder whether the recently adopted EU Artificial Intelligence Act (AIA) [^13] EU Regulation 2024/1689, 12.7.2025. provides an answer on how to address the memorization of personal data in LLMs. The AIA aims to provide a legal framework for the development, deployment, and use of human-centered and trustworthy artificial intelligence (AI) [^14] Article 1(1) AIA.. It establishes legal requirements for specific types of AI systems and general-purpose AI models that must be fulfilled before they are placed on the market, inter alia, distributed or used in the EU in the course of a (commercial) activity. [^15] Article 3(10) AIA. Without going into detail, it is important to note that the AI Act applies without prejudice to the GDPR. [^16] Article 2(7) AIA. Therefore, both legal frameworks could apply simultaneously. However, it remains questionable whether the AI Act would be relevant to our topic and to ML researchers since the AI Act does not apply to AI systems or AI models, including their output, that are developed and put into service solely for the purpose of scientific research and development. [^17] Article 2(6) AIA. It is indeed questionable to what extent this legal exemption applies to ML researchers who publish their LLMs on public platforms. Therefore, the issue of memorization of personal data remains relevant only under the GDPR.
From our point of view, a combination of technical and compliance measures will be the favorable approach to protect personal data. Additionally, awareness in the ML community about challenges and solutions is beneficial.
6 Summary
It is argued that LLMs do not contain personal data, and even if they did, extending data protection rights and obligations to LLMs would create unmanageable demands, particularly regarding rights to access and to be forgotten. Indeed, implementing these rights is challenging due to the nature and operational complexity of LLMs. Current research practices may be pushing the GDPR’s goal of technology neutrality to its limits.
However, practical difficulties in enforcing legal consequences do not negate the legal applicability. Reducing the scope of the GDPR based on unforeseen or conflicting interests would contradict its comprehensive protective purpose and intended technology neutrality. This could lead to significant legal uncertainty and potential liability for the ML community. Instead, solutions should focus on adapting legal consequences and developing new technical approaches for GDPR compliance.
If the ML community acknowledges the legal implications of LLMs as personal data, it can better engage with policymakers and other stakeholders to influence legislation. While a political shift in this direction is unlikely, the question of whether the GDPR needs to be adjusted for new AI technologies has already been raised (Sartor et al., 2020; Mitrou, 2018). Additionally, it remains to be seen how courts will address this issue (see pending cases such as (NOBY, 4)).
Overall, we encourage the ML community to acknowledge GDPR-related challenges in the development and deployment of LLMs. Joint efforts between computer science and law are needed to tackle the challenges.
Acknowledgments
We thank Finn Fassbender and Thomas Klein for their helpful feedback on this paper.
All authors acknowledge support from the Carl Zeiss Foundation. Moreover, Michèle Finck and Kristof Meding are members of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.